OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. d + e = 15 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (10, −5) (5, 10) (10, 5)

OpenStudy (carolinar7):

ok do you know what a system of equations is

OpenStudy (anonymous):

no

OpenStudy (carolinar7):

A system of equations is a collection of two or more equations with a same set of unknowns. In solving a system of equations, we try to find values for each of the unknowns that will satisfy every equation in the system. The equations in the system can be linear or non-linear.

OpenStudy (carolinar7):

Now that you know this would you rather solve it by elimination or substitution

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I solve the second equation with respect to e, and I find: \[\Large e = d - 5\]

OpenStudy (carolinar7):

Now substitue that into the other equation

OpenStudy (michele_laino):

what do you get? @lil_chukie

OpenStudy (carolinar7):

d+(d-5)=15

OpenStudy (michele_laino):

please continue @lil_chukie

OpenStudy (anonymous):

idk how to do this

OpenStudy (anonymous):

@Michele_Laino whats the answer

OpenStudy (michele_laino):

we have: 2d-5=15 then 2d=20 and d=20/2=...?